The manager of a store that specializes in selling tea decides to experiment with a new blend. She will mix some Earl Grey tea that sells for $4 per pound with some Orange Pekoe tea that sells for $2 per pound to get 200 pounds of the new blend. The selling price of the new blend is to be $2.50 per​ pound, and there is to be no difference in revenue from selling the new blend versus selling the other types. How many pounds of the Earl Grey tea and Orange Pekoe tea are​ required?
Answer by addingup(3677)

quantity of EG: x
quantity of OP: 200-x
--------------------
4x+2(200-x) = 2.50(200)
4x+400-2x = 500
2x = 100
:
John